OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively. an = 3 • (-4)^n + 1 an = 3 • 4^n - 1 an = 3 • (-4)^n - 1 an = 3 • 4^n

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ a_5=a_2 \times r^{3-1} }\)

OpenStudy (solomonzelman):

first find the common ratio.

OpenStudy (anonymous):

whats r??

OpenStudy (anonymous):

786=-12*r^3-1

OpenStudy (solomonzelman):

I think I was wrong

OpenStudy (anonymous):

so 786/-12=r^2

OpenStudy (anonymous):

oh ok

OpenStudy (solomonzelman):

\(\Large\color{black}{ a_5=a_2 \times r^{5-2} }\) \(\Large\color{black}{ a_5=a_2 \times r^{3} }\) \(\Large\color{black}{ 786=(-12) \times r^{3} }\)

OpenStudy (solomonzelman):

(this is also logical, because otherwise you get an imaginary ratio)

OpenStudy (anonymous):

ok so 768/-12=r^3 -64=r^3

OpenStudy (anonymous):

-4=r

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

So the first term \(\Large\color{black}{ a_1 }\) is equal to ? (Knowing that the 2nd term is `-12` , and ` r=-4 ` )

OpenStudy (anonymous):

3

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

Yes.

OpenStudy (anonymous):

so what do i do from there?

OpenStudy (solomonzelman):

The recursive formula (for the nth term) is like this, \(\Large\color{black}{ a_n=a_1 \times r^{(n-1)} }\) all you have to do is to `fill in your numbers`

OpenStudy (solomonzelman):

you found that \(\Large\color{blue}{ a_1=3 ~~~~~\rm{and}~~~~~r=-4}\)

OpenStudy (anonymous):

so its the third answer?

OpenStudy (solomonzelman):

yes, assuming that it is an = 3 • (-4)^\(\normalsize\color{red}{ (}\)n-1\(\normalsize\color{red}{ )}\) see what I am adding? It's important.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (solomonzelman):

Anytime! This was one of my favorite topics in school. Take care:)